Wednesday, 19 August

Other Events

12:15  TU Main Building, “Lichthof”

Terror and Exile Opening of an exhibition, organized by the German Mathematical Society (DMV), about the persecution and expulsion of mathematicians from Berlin between 1933 and 1945.

19:30  Auditorium Maximum H 105

Andrew Wiles, Princeton University, Princeton, USA Twenty Years of Number Theory Chair: Don Zagier, Max-Planck-Institut für Mathematik, Bonn, Germany

ICM’98 – Program  1